# Geometric, topological and dynamical properties of conformally symplectic systems, normally hyperbolic invariant manifolds, and scattering maps

### v1

@@ 9. Formulas for the primitive functions of wave maps and scattering map when $f$ is exact
When $f$ is exact conformally symplectic, in this section we obtain formulas for the primitive functions of the wave maps and the scattering map in terms of the primitive function of $f$. The variational formulation for conformally symplectic systems is given in (9.21). The formulas (9.6) (9.7) and (9.9) provide a link with the calculus of variations for conformally symplectic systems. In the symplectic case, [Ang93, Lom97, AB98] develop variational descriptions of heteroclinic connections. Such formulas have been used in numerical calculations of orbits homoclinic (or heteroclinic) to periodic orbits [Tab95] of twist maps.
@@
Given $x\in W^{s,\mathrm{loc}}_{\Lambda}$, when $\sigma$ is chosen to be the path $\gamma_{+}(\cdot;x,\Omega_{+}(x))\subset W^{s,\mathrm{loc}}_{\Omega_{+}(x)}$ given in (8.4), and denoting by $\gamma^{N}_{+}:=f^{N}(\gamma_{+}(\cdot;x,\Omega_{+}(x))=\gamma_{+}(\cdot;f^{N}(x),\Omega_{+}(f^{N}(x)))$, using the formula (8.23) for the primitive function $P^{+}_{\alpha}$ (and, similarly, in the analogous formula for $P^{-}_{\alpha}$), we obtain:
@@
However, even if $\omega$ is uniformly bounded, $\alpha$ may be unbounded (see Section 9.4 for lower bounds for all forms). If $f^{N}(x)$ escapes to infinity, it could happen that $\|\alpha_{f^{N}(x)}\|$ grows so fast that it overtakes the decrease of the length of $\gamma^{N}_{\pm}$.
@@
In Section 9.2 we will show that, if there are orbits that escape to infinity, there is always a gauge transformation that makes (9.6) and (9.7) divergent.
@@
This is a direct consequence of [Spi65, Theorem 3-11].
@@
The purpose of this section is to show that, if $f_{\mid\Lambda}$ has orbits that escape to infinity101010the orbit $f^{n}(y)$ escapes to infinity if every compact $K\subset\Lambda$ contains only finitely many terms of the subsequence, there is always a function $G$ (in fact, plenty of them) such that the series (9.6), which give the primitive $P_{\alpha}^{+}$ of the wave map $\Omega^{+}$, corresponding to the modified $1$-form $\alpha+dG$,
@@ 9.3. Variational interpretation of the iterative formulas for primitive functions of scattering map
Let $T^{*}Q$ be a cotangent bundle of a manifold – the standard example of a symplectic manifold – with the canonical $1$-form defined in coordinates $(p,q)$ as $\alpha_{0}=pdq$  [^11] A geometrically natural definition of $\alpha_{0}$ is standard, see, for example: [AM78, Proposition 3.2,11 p. 180].. Let $f$ be a mapping on $T^{*}Q$ that is homotopic to the identity, exact conformally symplectic ($f^{*}\alpha-\eta\alpha=dP^{f}_{\alpha}$ for some function $P^{f}_{\alpha}$ on $T^{*}Q$), and satisfies a twist condition, 121212 The twist condition is clearly non-generic – it is not verified by the identity– but it is verified by the geodesic flow of a compact manifold at time $t>0$ for small enough $t$ [Gol94]. meaning that if $f(p,q)=(p^{\prime},q^{\prime})$ then $(q,q^{\prime})$ gives a system of coordinates on $Q\times Q$. The primitive function $P^{f}_{\alpha}$ has the interpretation of an action and that the orbits of $f$ are critical points of the formal action
@@
where $x=(x_{n})_{n}=(f^{n}(x))_{n}$. See [Har00] for the symplectic case $\eta=1$.
@@
The variational principle in (9.21), often called discounted variational principle, appears naturally in finance. Each $x_{n}$ is a transaction at time $n$ and $P^{f}_{\alpha}(x_{n})$ is the cost of the transaction in currency. If there is constant inflation, to evaluate the cost of a strategy, it is natural to add the costs at different times by reducing them to a common time [Ben88]. The conversion of the cost $P^{f}_{\alpha}(x_{n})$ to currency at $n=0$ is $\eta^{-n}P^{f}_{\alpha}(x_{n})$.
The variational principles (9.21) also appear in control theory under the name finite horizon approximation [KKR17].
@@
The variational approach to homoclinic orbits has also been used as a numerical tool for symplectic systems [Tab95] [Tab95, MMS89]. Using that (9.4), (9.5) can be computed by other methods and considered as boundary terms, it seems that the method could be adapted to conformally symplectic systems.
@@ Appendix A Summary of the theory on properties of NHIMs
We consider the setting from Section 3.1 without assuming a conformal structure and that $f$ is conformally symplectic.
@@
Note that the mapping $(f,g)\rightarrow f\circ g$ is not continuous from $C^{0}\times C^{0}\rightarrow C^{0}$ unless $f$ is uniformly continuous. If one does not use fractional regularities for $f$, the only way to obtain uniform continuity is to assume $f\in C^{1}$. This may lead to extra losses of regularity in the conclusions. One possibility used in several references is to include the uniform continuity of the highest derivative in the definition of $C^{r}$ but note that, when the domain is unbounded, the uniform continuity is not preserved under uniform limits, so that the space thus defined is not a Banach space.
@@
The fact that if we fix $y\in W^{s,\textrm{loc}}_{\Lambda}$, $x\in\Lambda$ is determined uniquely, as claimed in (II.vi) of Theorem A.1, is crucial for this paper, since the projection from $y$ to $x$ – defined by $\Omega_{+}$ – is an important ingredient. This property requires assumption (U2). [Eld12, Example 2.9] (See Figure 2) provides an example of NHIM where were $W^{s}_{x}\cup W^{s}_{x^{\prime}}\neq\emptyset$ for $x\neq x^{\prime}$ so that the partition of the stable manifold into strong stable leaves is not a foliation.
@@
In the study of NHIM, it is natural to consider two distances on $\Lambda$: $d_{\Lambda}(x,x^{\prime})$ is the shortest length of paths in $\Lambda$ joining $x,x^{\prime}$, and $d_{M}(x,x^{\prime})$ is the shortest length of paths in $M$. Clearly, $d_{M}(x,x^{\prime})\leq d_{\Lambda}(x,x^{\prime})$. The distance that enters in the definition of $W^{s}_{x}$ is $d_{M}$. It is easy to see that for $\varepsilon>0$, $x,x^{\prime}\in\Lambda$, $d_{\Lambda}(f^{n}(x),f^{n}(x^{\prime}))\leq C(\lambda_{+}-\varepsilon)^{n},n>0$ implies $x=x^{\prime}$. On the other, hand, if $\Lambda$ folds into itself (as in [Eld12, Example 2.9]) it is possible that for two different points $x,x^{\prime}\in\Lambda$, $d_{M}(f^{n}(x),f^{n}(x^{\prime}))\leq C\lambda_{+}^{n}$.
@@
It is standard that $A$ can be interpreted either as a diffeomorphism of ${\mathbb{T}}^{d}$ (taking the action mod $1$) or as a linear map on ${\mathbb{Z}}^{d}$ (acting on the Fourier coefficients) [KH95].
@@
Analyzing these solutions reveals that, in many cases, they possess only a finite number of continuous derivatives. We will see that the analysis is very similar to the analysis of the classical Weierstrass example 151515 [Har16] considers the harder problem of no derivative of Weierstrass function at any point. .
A version of the argument close to the one here appears in [dlL92, Section 6.2]. Similar arguments appears often in hyperbolic systems.
@@
There are many proofs of Theorem A.1, some of them have more assumptions. For the purposes of this paper, an efficient proof of the local stable/unstable manifolds is in [Pes04, p. 33–38]. This proof is based on constructing first the embeddings giving the local stable/unstable manifolds by studying the functional equations satisfied via a fixed point argument. Note that this proof requires uniformity assumptions for the map or the derivative in a neighborhood of $\Lambda$.
@@
Nevertheless, [Eld12, Example 3.8] (Illustrated here in Figure 2 ) shows that to get (II).(vi)(II).(vi),$(II).(vii)$, $(III)$ and all the other properties of $W^{s,\textrm{loc}}_{x}$ in Theorem A.1 one needs the uniformity assumption (U2). For the purposes of this paper, the fact that $W^{s}_{\Lambda}$ is foliated by $W^{s}_{x}$ is crucial since it is what we use to define $\Omega_{+}$ and the scattering map.
@@
The paper [Fen74] establishes regularity based on the study of the numbers $\alpha$ for which
@@
which is sharper than taking the supremum on all factors. as is done in [HPS77].
Conditions similar to (A.4) are used in the invariant cone approach to regularity of invariant manifolds [CZ15].
@@ B.2.1. A sharp result on perturbations of products of a sequence of operators (cocycles)
Using the system of coordinates in $\mathcal{O}_{\rho}$ we can use connectors (see (2.1)) to identify $E^{s}_{x}$ with $T_{y}W^{s}_{x}$. A geometrically natural way to identify these spaces is to take $S_{x}^{y}=(D\exp_{x})(\exp^{-1}_{x}(y))$ where $\exp$ denotes the geometric exponential mapping along the manifold $W^{s}_{x}$ (intuitively, if $y=\exp_{x}(v)$, $S_{x}^{y}w=\frac{d}{dt}\exp_{x}(v+tw)|_{t=0}$). We note that the norm of these operators is bounded and in $\mathcal{O}_{\rho}$ and the norm becomes close to $1$ if $\rho$ is small.
@@ Appendix C Rates of convergence of homoclinic channels to a NHIM
The Lemma C.1 is closely related to results in the literature such as the graph transform, the inclination lemma – a.k.a. $\lambda$-lemma, and even closer, to fiber contraction lemma, or the more general cone conditions, which have many variants. In our case, the result is easier since we already know the fixed point of the contraction. Unfortunately, many of the versions in the literature are only qualitative or involve extra assumptions (e.g. [HPS77, p. 35] assumes compactness).
The tangent functor trick [AR67, HP70] – which we explain later – shows that the case $j=1$ implies the results for other $j$ in Lemma C.1. The case $j=1$ we present, basically goes back to [Had98] (translated to English in [Has17]).
@@
For future reference, we compute some explicit and elementary expressions for the derivatives of iterations of $f$ (similarly to [DdlLS08, Proposition 11]).
@@
Following [Had01, Has17], we consider $(\Delta,(D\sigma(x))\Delta)$ the representation of a tangent vector of $\Gamma$ at $(x,\sigma(x))$. When $\Delta$ ranges over all the possible values, $(\Delta,(D\sigma(x))\Delta)$ ranges over the graph of $D\sigma(x)$, the tangent space of $\Gamma$.
@@
For this, we use the ‘tangent functor trick’ [AR67, HP70]. Note that given $f:M\rightarrow N$, $g:N\rightarrow P$, differentiable maps among manifolds, defining $Tf:TM\rightarrow TN$, $Tg:TN\rightarrow TP$ by $Tf(x,v)=(f(x),Df(x)v)$, we have:
@@ D.1. Partially conformally symplectic systems
Systems of the form (D.1) appear in mechanics when we consider models of $L$ particles interacting by a Hamiltonian, where each particle is subject to a friction proportional to its velocity. See [CCdlL13, Remark 3].
@@
  • The spin-spin model [^16] The spin-spin model is a time dependent flow. The time advance maps are of the form (D.1). describing two rotating bodies, each with its own tidal friction [Mis21, eq. (42)].
  • Networks of oscillators with Hamiltonian coupling among nearest neighbors, where one node experiences dissipative effects [EW20, EW18, CEW17, DFKY25]. These are models of the form (D.1) in which $\eta_{i}=1\textrm{ for }i\neq 1$, $\eta_{1}<1$. There are invariant manifolds that affect the transfer of energy from the conservative modes to the dissipative modes.
  • Models of planets with a two-layer structure subject to viscous and tidal friction, each of them with different friction coefficients. See [PSV24].
@@
which is a system of uncoupled conformally symplectic oscillators with uniform dissipation. Other models with strong dissipation $\frac{d}{dt}u_{xx}$ instead of $u_{t}$ have also been considered. Similar considerations apply to other Hamiltonian Partial Differential equations and infinite dimensional coupled systems. For example, [EW20, EW18, CEW17] explore analogies with locally dissipative variants of the nonlinear Schrödinger equation.
Making rigorous sense of geometric properties of NHIMs in PDEs seems an interesting problem, but it is tractable for finite dimensional manifolds [BLZ08].
One should note that in infinite dimensional models, there may be dissipation of energy even in Hamiltonian systems [CG22, CKK23].

### v2

@@ 9. Formulas for the primitive functions of wave maps and scattering map when $f$ is exact
When $f$ is exact conformally symplectic, in this section we obtain formulas for the primitive functions of the wave maps and the scattering map in terms of the primitive function of $f$. The variational formulation for conformally symplectic systems is given in (9.21). The formulas (9.6) (9.7) and (9.9) provide a link with the calculus of variations for conformally symplectic systems. In the symplectic case, [Angenent93, Lomeli97, AmbrosettiB98] develop variational descriptions of heteroclinic connections. Such formulas have been used in numerical calculations of orbits homoclinic (or heteroclinic) to periodic orbits [Tabacman95] of twist maps.
@@
Given $x\in W^{s,\mathrm{loc}}_{\Lambda}$, when $\sigma$ is chosen to be the path $\gamma_{+}(\cdot;x,\Omega_{+}(x))\subset W^{s,\mathrm{loc}}_{\Omega_{+}(x)}$ given in (8.4), and denoting by
$\gamma^{N}_{+}:=f^{N}(\gamma_{+}(\cdot;x,\Omega_{+}(x))=\gamma_{+}(\cdot;f^{N}(x),\Omega_{+}(f^{N}(x))),$
using the formula (8.22) for the primitive function $P^{+}_{\alpha}$ (and, similarly, in the analogous formula for $P^{-}_{\alpha}$), we obtain:
@@
However, even if $\omega$ is uniformly bounded, $\alpha$ may be unbounded (see Section 9.4 for lower bounds for all forms). If $f^{N}(x)$ escapes to infinity, it could happen that $\|\alpha_{f^{N}(x)}\|$ grows so fast that it overtakes the decrease of the length of $\gamma^{N}_{\pm}$ (and the powers of $\eta$).
@@
In Section 9.2 we will show that, if there are orbits that escape to infinity, there is always a gauge that makes (9.6) and (9.7) divergent.
@@
This is a direct consequence of [Spivak, Theorem 3-11].
@@
The purpose of this section is to show that, if $f_{\mid\Lambda}$ has orbits that escape to infinity121212the orbit $f^{n}(y)$ escapes to infinity if every compact $K\subset\Lambda$ contains only finitely many points of the orbit, there is always a function $G$ (in fact, plenty of them) such that the series (9.6), which give the primitive $P_{\alpha}^{+}$ of the wave map $\Omega^{+}$, corresponding to the modified $1$-form $\alpha+dG$,
@@ 9.3. Variational interpretation of the iterative formulas for primitive functions of scattering map
Let $T^{*}Q$ be a cotangent bundle of a manifold – the standard example of a symplectic manifold – with the canonical $1$-form defined in coordinates $(p,q)$ as $\alpha_{0}=pdq$  [^13] A geometrically natural definition of $\alpha_{0}$ is standard, see, for example: [AbrahamM78, Proposition 3.2,11 p. 180].. Let $f$ be a mapping on $T^{*}Q$ that is homotopic to the identity, exact conformally symplectic ($f^{*}\alpha-\eta\alpha=dP^{f}_{\alpha}$ for some function $P^{f}_{\alpha}$ on $T^{*}Q$), and satisfies a twist condition,  [^14] The twist condition is clearly non-generic – it is not verified by the identity– but it is verified by the geodesic flow of a compact manifold at time $t>0$ for small enough $t$ [Gole94]. meaning that if $f(p,q)=(p^{\prime},q^{\prime})$ then $(q,q^{\prime})$ gives a system of coordinates on $Q\times Q$. The primitive function $P^{f}_{\alpha}$ has the interpretation of an action and that the orbits of $f$ are critical points of the formal action
@@
where $x=(x_{n})_{n}=(f^{n}(x))_{n}$. See [Haro00] for the symplectic case $\eta=1$.
@@
The variational principle in (9.21), often called discounted variational principle, appears naturally in finance. Each $x_{n}$ is a transaction at time $n$ and $P^{f}_{\alpha}(x_{n})$ is the cost of the transaction in currency. If there is constant inflation, to evaluate the cost of a strategy, it is natural to add the costs at different times by reducing them to a common time [Bensoussan88]. The conversion of the cost $P^{f}_{\alpha}(x_{n})$ to currency at $n=0$ is $\eta^{-n}P^{f}_{\alpha}(x_{n})$.
The variational principles (9.21) also appear in control theory under the name finite horizon approximation [KaliseKR17].
@@
The variational approach to homoclinic orbits has also been used as a numerical tool for symplectic systems [Tabacman95, MacKayMS89]. Using that (9.4), (9.5) can be computed by other methods and considered as boundary terms, it seems that the method could be adapted to conformally symplectic systems.
@@ Appendix A Summary of the theory on properties of NHIMs
We consider the setting from Section 3.1 without assuming a symplectic structure on the manifold and that the map is conformally symplectic.
@@
Note that the mapping $(f,g)\rightarrow f\circ g$ is not continuous from $C^{0}\times C^{0}\rightarrow C^{0}$ unless $f$ is uniformly continuous. If one does not use fractional regularities for $f$, the only way to obtain uniform continuity for $f$ is to assume $f\in C^{1}$. This may lead to extra losses of regularity in the conclusions. One possibility used in several references is to include the uniform continuity of the highest derivative in the definition of $C^{r}$ but note that, when the domain is unbounded, the uniform continuity is not preserved under uniform limits, so that the space thus defined is not a Banach space.
@@
The fact that if we fix $y\in W^{s,\textrm{loc}}_{\Lambda}$, $x\in\Lambda$, is determined uniquely, as claimed in (II.vi) of Theorem A.1, is crucial for this paper, since the projection from $y$ to $x$ – defined by $\Omega_{+}$ – is an important ingredient. This property requires assumption (U2). [Eldering12, Example 2.9] (See Figure 2) provides an example of NHIM for which $W^{s}_{x}=W^{s}_{x^{\prime}}$ for some $x\neq x^{\prime}$. Therefore, the decomposition of $W^{s}_{\Lambda}$ into strong stable leaves $W^{s,loc}_{x}$ in (2.15) is not a disjoint union  [^17] Since the relation given by $y\approx\tilde{y}\iff d(f^{n}(y),f^{n}(\tilde{y}))\leq C_{y,\tilde{y}}\lambda_{+}^{n},n\geq 0$ is an equivalence relation, the equivalence classes under this relation are leaves of a local foliation. Nevertheless, in [Eldering12, Example 2.9] a leaf can contain two points in $\Lambda$, so that the leaves cannot be labelled by their point of intersection with $\Lambda$. For our purposes, this prevents us from defining $\Omega_{+}$. .
@@
In the study of NHIM, it is natural to consider two distances on $\Lambda$: $d_{\Lambda}(x,x^{\prime})$ is the shortest length of paths in $\Lambda$ joining $x,x^{\prime}$, and $d_{M}(x,x^{\prime})$ is the shortest length of paths in $M$. Clearly, $d_{M}(x,x^{\prime})\leq d_{\Lambda}(x,x^{\prime})$. The distance that enters in the definition of $W^{s}_{x}$ is $d_{M}$. It is easy to see that for $\varepsilon>0$, $x,x^{\prime}\in\Lambda$, $d_{\Lambda}(f^{n}(x),f^{n}(x^{\prime}))\leq C(\lambda_{+}-\varepsilon)^{n},n>0$ implies $x=x^{\prime}$. On the other, hand, if $\Lambda$ folds into itself (as in [Eldering12, Example 2.9]) it is possible that for two different points $x,x^{\prime}\in\Lambda$, $d_{M}(f^{n}(x),f^{n}(x^{\prime}))\leq C\lambda_{+}^{n}$.
@@
It is standard that $A$ can be interpreted either as a diffeomorphism of ${\mathbb{T}}^{d}$ (taking the action mod $1$) or as a linear map on ${\mathbb{Z}}^{d}$ (acting on the Fourier coefficients) [KatokH95].
@@
Analyzing these solutions reveals that, in many cases, they possess only a finite number of continuous derivatives. We will see that the analysis is very similar to the analysis of the classical Weierstrass example 181818[Hardy16] considers the harder problem of no derivative of Weierstrass function at any point. .
A version of the argument close to the one here appears in [Llave92, Section 6.2]. Similar arguments appears often in hyperbolic systems.
@@
There are many proofs of Theorem A.1, some of them have more assumptions. For the purposes of this paper, an efficient proof of the local stable/unstable manifolds is in [Pesin, p. 33–38]. This proof is based on constructing first the embeddings giving the local stable/unstable manifolds by studying the functional equations satisfied via a fixed point argument. Note that this proof requires uniformity assumptions for the map or the derivative in a neighborhood of $\Lambda$.
@@
Nevertheless, [Eldering12, Example 3.8] (Illustrated here in Figure 2 ) shows that to get (II).(vi)(II).(vi),$(II).(vii)$, $(III)$ and all the other properties of $W^{s,\textrm{loc}}_{x}$ in Theorem A.1 one needs the uniformity assumption (U2). For the purposes of this paper, the fact that $W^{s}_{\Lambda}$ is foliated by $W^{s}_{x}$ is crucial since it is what we use to define $\Omega_{+}$ and the scattering map.
@@
The paper [Fenichel74] establishes regularity based on the study of the numbers $\alpha$ for which
@@
which is sharper than taking the supremum on all factors. as is done in [HPS].
Conditions similar to (A.4) are used in the invariant cone approach to regularity of invariant manifolds [CapinskiZ15].
@@ B.2.1. A sharp result on perturbations of products of a sequence of operators (cocycles)
Using the system of coordinates in $\mathcal{O}_{\rho}$ we can use connectors (see (2.1)) to identify $E^{s}_{x}$ with $T_{y}W^{s}_{x}$. A geometrically natural way to identify these spaces is to take $S_{x}^{y}=(D\exp_{x})(\exp^{-1}_{x}(y))$. We note that the norm of these operators is bounded and in $\mathcal{O}_{\rho}$ and the norm becomes close to $1$ if $\rho$ is small.
@@ Appendix C Rates of convergence of homoclinic channels to a NHIM
The Lemma C.1 is closely related to results in the literature such as the graph transform, the inclination lemma – a.k.a. $\lambda$-lemma, and even closer, to fiber contraction lemma, or the more general cone conditions, which have many variants. In our case, the result is easier since we already know the fixed point of the contraction. Unfortunately, many of the versions in the literature are only qualitative or involve extra assumptions (e.g. [HPS, p. 35] assumes compactness).
The tangent functor trick [AbrahamR67, HirschP70] – which we explain later – shows that the case $j=1$ implies the results for other $j$ in Lemma C.1. The case $j=1$ we present, basically goes back to [Hadamard98] (translated to English in [Hasselblatt17]).
@@
For future reference, we compute some explicit and elementary expressions for the derivatives of iterations of $f$ (similarly to [DLS08, Proposition 15]).
@@
Following [Hadamard01, Hasselblatt17], we consider $(\Delta,(D\sigma(x))\Delta)$ the representation of a tangent vector of $\Gamma$ at $(x,\sigma(x))$. When $\Delta$ ranges over all the possible values, $(\Delta,(D\sigma(x))\Delta)$ ranges over the graph of $D\sigma(x)$, the tangent space of $\Gamma$.
@@
For this, we use the ‘tangent functor trick’ [AbrahamR67, HirschP70]. Note that given $f:M\rightarrow N$, $g:N\rightarrow P$, differentiable maps among manifolds, defining $Tf:TM\rightarrow TN$, $Tg:TN\rightarrow TP$ by $Tf(x,v)=(f(x),Df(x)v)$, we have:
@@ D.1. Partially conformally symplectic systems
Systems of the form (D.1) appear in mechanics when we consider models of $L$ particles interacting by a Hamiltonian, where each particle is subject to a friction proportional to its velocity. See [CallejaCL13, Remark 3].
@@
  • The spin-spin model [^19] The spin-spin model is a time dependent flow. The time advance maps are of the form (D.1). describing two rotating bodies, each with its own tidal friction [Misquero21, eq. (42)].
  • Networks of oscillators with Hamiltonian coupling among nearest neighbors, where one node experiences dissipative effects [EckmannW20, EckmanW18, CuneoEW17, DolgopyatFKY2025]. These are models of the form (D.1) in which $\eta_{i}=1\textrm{ for }i\neq 1$, $\eta_{1}<1$. There are invariant manifolds that affect the transfer of energy from the conservative modes to the dissipative modes.
  • Models of planets with a two-layer structure subject to viscous and tidal friction, each of them with different friction coefficients. See [PinzariSV24].
@@
which is a system of uncoupled conformally symplectic oscillators with uniform dissipation. Other models with strong dissipation $\frac{d}{dt}u_{xx}$ instead of $u_{t}$ have also been considered. Similar considerations apply to other Hamiltonian Partial Differential equations and infinite dimensional coupled systems. For example, [EckmannW20, EckmanW18, CuneoEW17] explore analogies with locally dissipative variants of the nonlinear Schrödinger equation.
Making rigorous sense of geometric properties of NHIMs in PDEs seems an interesting problem, but it is tractable for finite NHIMs [BatesLZ08].
One should also note that in infinite dimensional models, there may be dissipation of energy even in Hamiltonian systems [CorsiG22, ComechKK23].